OpenStudy (anonymous):

use cramer's rule/ determinants to solve 8x-3y=5 2x+9y=-2

OpenStudy (anonymous):

Are you solving for system of equation?

OpenStudy (anonymous):

I can help you if you are.

OpenStudy (anonymous):

Express the system using matrices: \[\begin{bmatrix}8&-3\\2&9\end{bmatrix}\begin{bmatrix}x\\y\end{bmatrix}=\begin{bmatrix}\color{red}5\\\color{red}{-2}\end{bmatrix}\] Cramer's rule can be used if the coefficient matrix has non-zero determinant: \[\begin{vmatrix}8&-3\\2&9\end{vmatrix}=72-(-6)=78\not=0\] Conditions are met, so we can use Cramer's to solve directly for \(x\) and \(y\): \[x=\frac{\begin{vmatrix}\color{red}5&-3\\\color{red}{-2}&9\end{vmatrix}}{\begin{vmatrix}8&-3\\2&9\end{vmatrix}}~~\text{and}~~y=\frac{\begin{vmatrix}8&\color{red}5\\2&\color{red}{-2}\end{vmatrix}}{\begin{vmatrix}8&-3\\2&9\end{vmatrix}}\]